# MathEDU: Towards Adaptive Feedback for Student Mathematical Problem-Solving

### v1

@@ Appendix I Mathematical Problem-Solving Ability of the Model
Appendix L Analysis of Model-Generated Feedback
In Table 28, another example is presented where the question asks which orientation of a water tank placed in a crate measuring 8 by 12 by 14 feet would yield the largest volume. The tank is to be placed upright, and the student mistakenly used the smallest dimension divided by 2 as the radius, leading to an error. The problem requires consideration of different orientations of the crate, ultimately determining that placing the crate on the 12 by 14 face yields the maximum tank volume, with the correct radius being 6 feet.
Llama3 70B recognized that the student made an error but ended up performing the same calculation as the student, arriving at the same incorrect answer, which was unhelpful. Similarly, GPT-3.5 also failed to consider the varying tank volumes based on different crate orientations. It acknowledged the student’s mistake but provided the same incorrect answer, rendering it similarly unhelpful.
@@
[…]
@@
Appendix M Information About Use Of Ai Assistants
We utilized generative AI tools solely for grammar and language refinement. Following the use of this tool, the author(s) carefully reviewed and revised the content as necessary, assuming full responsibility for the final publication.

### v2

@@ Appendix F Evaluation Criteria for Generated Feedback Quality
o1-mini: The student correctly calculated the total amount of apples sold by adding the quantities from each hour: $10\,\text{kg}+2\,\text{kg}=12\,\text{kg}$ However, the question asks for the average amount of apples sold per hour, not the total. To find the average: 1. Total apples sold: $10\,\text{kg}+2\,\text{kg}=12\,\text{kg}$ 2. Total time: $2\,\text{hours}$ 3. Average per hour: $\frac{12\,\text{kg}}{2\,\text{hours}}=6\,\text{kg/hour}$ Advice: Always pay close attention to whether a question is asking for a total, an average, or another specific calculation. In this case, dividing the total by the number of hours provides the correct average.
@@
Appendix G Case Studies of Model-Generated Feedback
Table 19 presents another example where the question asks which orientation of a water tank placed in a crate measuring 8 by 12 by 14 feet would yield the largest volume. The tank is to be placed upright, and the student mistakenly used the smallest dimension divided by 2 as the radius, leading to an error. The problem requires consideration of different orientations of the crate, ultimately determining that placing the crate on the 12 by 14 face yields the maximum tank volume, with the correct radius being 6 feet. Llama3 70B recognized that the student made an error but ended up performing the same calculation as the student, arriving at the same incorrect answer, which was unhelpful. Similarly, GPT-3.5 also failed to consider the varying tank volumes based on different crate orientations. It acknowledged the student’s mistake but provided the same incorrect answer, rendering it similarly unhelpful.
Appendix H GenAI Usage Disclosure
We utilized generative AI tools solely for grammar and language refinement. All content was reviewed and edited by the author(s), who take full responsibility for the final manuscript.